# Semantic-Guided Two-Stage GAN for Face Inpainting with Hybrid Perceptual Encoding

### v1

@@ 1 Introduction
The task of filling missing pixels of an image with meaningful content, often referred to as image inpainting or completion, is an important task in computer vision. It has seen significant progress with deep learning advances. [Pathak_2016_CVPR, Yu_2018_CVPR]. However facial image inpainting creates unique challenges due to its highly structured nature of facial geometry, the need for semantic coherence across facial components (eyes, nose, mouth), and human sensitivity to facial distortions [Liu_2018_ECCV, Nazeri_2019_ICCV]. Current state-of-the-art methods can be categorized into three paradigms: attention-based approaches [Yu_2018_CVPR, Zeng_2019_CVPR], two-stage coarse-to-fine refinement [Liu_2018_ECCV, Xiong_2019_CVPR], and transformer-based architectures [dosovitskiy2020image, Esser_2021_CVPR]. While these methods have achieved impressive results, they face several persistent limitations:
Semantic Inconsistency: Methods that directly predict RGB pixels tend to result in an image that violates constraints on facial structure, often featuring unrealistic characteristics, such as misaligned eyes or distorted facial boundaries [Liu_2018_ECCV, yan2018shift]
Texture Blurriness: Methods based on an optimization process, having only either of $\ell_{1}$ or $\ell_{2}$ reconstruction losses, tend to produce over-smooth results devoid of high-frequency details[Pathak_2016_CVPR, iizuka2017globally], leading to the usual ”blurry” appearance that degrades perceptual quality.
Boundary Artifacts: Insufficient attention in mask boundaries, causes noticeable lines or color mismatches between the inpainted and known regions of the image [Yu_2019_ICCV, Liu_2019_ICCV]. Destroying the overall smoothness and makes the result look less natural.
Limited Diversity: Models that always produce the same result struggle to show the many possible ways an image could be completed, producing less realistic images [zhao2021large, Zheng_2019_CVPR].
@@ 2.1 Classical Image Inpainting
Early inpainting methods relied on patch-based synthesis [barnes2009patchmatch, criminisi2004region] and diffusion-based propagation [bertalmio2000v, telea2004image]. Those approaches worked well for textures and simple patterns but failed on complex structures like faces due to their inability to understand high-level semantics features.
@@ 2.2 Deep Learning-Based Inpainting
Context Encoders: Pathak_2016_CVPR [Pathak_2016_CVPR] pioneered deep inpainting using an encoder-decoder architecture trained with adversarial and reconstruction losses. However, their method produced blurry results due to limited perceptual constraints.
Coarse-to-Fine Refinement: Liu_2018_ECCV [Liu_2018_ECCV]introduced a two-stage approach where a coarse network predicts initial results refined by a second network. Our work extends this paradigm but explicitly models semantic layouts rather than directly predicting RGB values in the coarse stage.
Attention Mechanisms: Yu_2018_CVPR [Yu_2018_CVPR]proposed contextual attention that copied features from known regions to missing areas. Zeng_2019_CVPR [Zeng_2019_CVPR] enhanced this with learnable bidirectional attention. We integrated multi-scale contextual attention but combined it with semantic guidance for better structural coherence.
@@ 2.3 Semantic-Guided Inpainting
As structure is important for inpainting tasks, several works used semantic guidance. yan2018shift [yan2018shift] used facial landmarks for guidance. Song_2018_ECCV [Song_2018_ECCV] used semantic parsing. Our approach generates probabilistic semantic layouts directly from masked inputs rather than requiring external semantic annotations.
@@ 2.4 Transformer-Based Methods
Recent works utilizes transformers for long-range dependency modeling. dosovitskiy2020image [dosovitskiy2020image] introduced Vision Transformers (ViT) for image recognition. Esser_2021_CVPR [Esser_2021_CVPR] applied transformers to image generation with VQGAN. Our hybrid CNN-Transformer encoder combines the local inductive biases of CNNs with the global reasoning of transformers, especially Vision Transformers.
@@ 2.5 Perceptual Quality Enhancement
johnson2016perceptual [johnson2016perceptual] introduced perceptual loss using pretrained VGG features. zhang2018unreasonable [zhang2018unreasonable] showed its effectiveness for various generation tasks. We used multi-scale perceptual loss to capture both low-level textures and high-level semantics.
@@ 2.6 GAN Training Stability
Training stability remains crucial for high-quality generation. gulrajani2017improved [gulrajani2017improved] proposed WGAN-GP with gradient penalty for stable training. karras2019style [karras2019style, karras2020analyzing] introduced progressive growing and style-based architectures. We adopted WGAN-GP with progressive loss scheduling for stable convergence.
@@ 3.3 Discriminator Design
Global Discriminator $D_{g}$ checks overall image realism using a standard CNN with spectral normalization [miyato2018spectral] .
Local Discriminator (PatchGAN) $D_{l}$ [isola2017image] assesses local texture realism by classifying overlapping patches, resulting in high-frequency detail generation.
@@ 3.4 Loss Functions
Multi-Scale Perceptual Loss: We extracted features from multiple VGG-19 layers $\phi_{l}$ [johnson2016perceptual]:
@@
WGAN-GP Adversarial Loss: We adopted Wasserstein GAN with gradient penalty [gulrajani2017improved] for stable training:
@@ 4.1 Experimental Setup
  • CelebA-HQ [karras2017progressive]: 30,000 high-quality facial images at 128×128 resolution
  • FFHQ [karras2019style]: 70,000 diverse faces at 128×128 resolution
  • Places2 [zhou2017places]: 1.8M scene images for generalization testing and experimentation.
@@
  • FID: Fréchet Inception Distance [heusel2017gans] (lower the better)
  • LPIPS: Learned perceptual image patch similarity [zhang2018unreasonable] (lower the better)
@@ 4.2 Comparison with State-of-the-Art
Our model is trained and evaluated on 128×128 CelebA images, whereas most state-of-the-art inpainting methods (e.g., DeepFill v2 [Yu_2019_ICCV], LaMa [Suvorov_2022_WACV], EdgeConnect [Nazeri_2019_ICCV], AOT‑GAN [zeng2022aggregated], MAT [Li_2022_CVPR]) and Structure matters [liu2024structure] are trained on 512×512 images. For this reason of resolution mismatch, direct quantitative comparisons using metrics such as PSNR, SSIM, LPIPS, and FID would be unfair and potentially misleading. To address this, we focussed on detailed qualitative comparisons, demonstrating the visual quality, texture consistency, and structural recovery of our method. Additionally, we perform ablation studies and reported quantitative metrics on our 128×128 validation set, highlighting the contribution of each component in our approach.

### v2

@@ 1 Introduction
The task of filling missing pixels of an image with meaningful content, often referred to as image inpainting or completion, is an important task in computer vision. It has seen significant progress with deep learning advances. [20, 26]. However facial image inpainting creates unique challenges due to its highly structured nature of facial geometry, the need for semantic coherence across facial components (eyes, nose, mouth), and human sensitivity to facial distortions [15, 19]. Current state-of-the-art methods can be categorized into three paradigms: attention-based approaches [26, 28], two-stage coarse-to-fine refinement [15, 24], and transformer-based architectures [4, 5]. While these methods have achieved impressive results, they face several persistent limitations:
Semantic Inconsistency: Methods that directly predict RGB pixels tend to result in an image that violates constraints on facial structure, often featuring unrealistic characteristics, such as misaligned eyes or distorted facial boundaries [15, 25]
Texture Blurriness: Methods based on an optimization process, having only either of $\ell_{1}$ or $\ell_{2}$ reconstruction losses, tend to produce over-smooth results devoid of high-frequency details[20, 8], leading to the usual ”blurry” appearance that degrades perceptual quality.
Boundary Artifacts: Insufficient attention in mask boundaries, causes noticeable lines or color mismatches between the inpainted and known regions of the image [27, 17]. Destroying the overall smoothness and makes the result look less natural.
Limited Diversity: Models that always produce the same result struggle to show the many possible ways an image could be completed, producing less realistic images [31, 32].
@@ 2.1 Classical Image Inpainting
Early inpainting methods relied on patch-based synthesis [1, 3] and diffusion-based propagation [2, 23]. Those approaches worked well for textures and simple patterns but failed on complex structures like faces due to their inability to understand high-level semantics features.
@@ 2.2 Deep Learning-Based Inpainting
Context Encoders: Pathak et al. [20] pioneered deep inpainting using an encoder-decoder architecture trained with adversarial and reconstruction losses. However, their method produced blurry results due to limited perceptual constraints.
Coarse-to-Fine Refinement: Liu et al. [15]introduced a two-stage approach where a coarse network predicts initial results refined by a second network. Our work extends this paradigm but explicitly models semantic layouts rather than directly predicting RGB values in the coarse stage.
Attention Mechanisms: Yu et al. [26]proposed contextual attention that copied features from known regions to missing areas. Zeng et al. [28] enhanced this with learnable bidirectional attention. We integrated multi-scale contextual attention but combined it with semantic guidance for better structural coherence.
@@ 2.3 Semantic-Guided Inpainting
As structure is important for inpainting tasks, several works used semantic guidance. Yan et al. [25] used facial landmarks for guidance. Song et al. [21] used semantic parsing. Our approach generates probabilistic semantic layouts directly from masked inputs rather than requiring external semantic annotations.
@@ 2.4 Transformer-Based Methods
Recent works utilizes transformers for long-range dependency modeling. Dosovitskiy [4] introduced Vision Transformers (ViT) for image recognition. Esser et al. [5] applied transformers to image generation with VQGAN. Our hybrid CNN-Transformer encoder combines the local inductive biases of CNNs with the global reasoning of transformers, especially Vision Transformers.
@@ 2.5 Perceptual Quality Enhancement
Johnson et al. [10] introduced perceptual loss using pretrained VGG features. Zhang et al. [30] showed its effectiveness for various generation tasks. We used multi-scale perceptual loss to capture both low-level textures and high-level semantics.
@@ 2.6 GAN Training Stability
Training stability remains crucial for high-quality generation. Gulrajani et al. [6] proposed WGAN-GP with gradient penalty for stable training. Karras et al. [12, 13] introduced progressive growing and style-based architectures. We adopted WGAN-GP with progressive loss scheduling for stable convergence.
@@ 3.3 Discriminator Design
Global Discriminator $D_{g}$ checks overall image realism using a standard CNN with spectral normalization [18] .
Local Discriminator (PatchGAN) $D_{l}$ [9] assesses local texture realism by classifying overlapping patches, resulting in high-frequency detail generation.
@@ 3.4 Loss Functions
Multi-Scale Perceptual Loss: We extracted features from multiple VGG-19 layers $\phi_{l}$ [10]:
@@
WGAN-GP Adversarial Loss: We adopted Wasserstein GAN with gradient penalty [6] for stable training:
@@ 4.1 Experimental Setup
  • CelebA-HQ [11]: 30,000 high-quality facial images at 128×128 resolution
  • FFHQ [12]: 70,000 diverse faces at 128×128 resolution
  • Places2 [33]: 1.8M scene images for generalization testing and experimentation.
@@
  • FID: Fréchet Inception Distance [7] (lower the better)
  • LPIPS: Learned perceptual image patch similarity [30] (lower the better)
@@ 4.2 Comparison with State-of-the-Art
Our model is trained and evaluated on 128×128 CelebA images, whereas most state-of-the-art inpainting methods (e.g., DeepFill v2 [27], LaMa [22], EdgeConnect [19], AOT‑GAN [29], MAT [14]) and Structure matters [16] are trained on 512×512 images. For this reason of resolution mismatch, direct quantitative comparisons using metrics such as PSNR, SSIM, LPIPS, and FID would be unfair and potentially misleading. To address this, we focussed on detailed qualitative comparisons, demonstrating the visual quality, texture consistency, and structural recovery of our method. Additionally, we perform ablation studies and reported quantitative metrics on our 128×128 validation set, highlighting the contribution of each component in our approach.